OpenStudy (anonymous):

How many grams of ethane gas (C2H6) are in a 12.7 liter sample at 1.6 atmospheres and 24°C? Show all work used to solve this problem.

sam (.sam.):

Use ideal gas law PV=nRT

sam (.sam.):

Find the moles first, \[n=\frac{PV}{RT}\] \[\Huge n=\frac{(1.6~atm)(12.7~L)}{(0.082\frac{atmL}{molK})(297.2K)}\] n=0.83mol ------------------------------------------------------------------ \[0.83~mol C_2H_{6} \times \frac{30~gC_2H_6}{1~molC_2H_6}=24.9gC_2H_6\]